### v1

@@ Abstract
Determinantal varieties—the sets of bounded-rank matrices or tensors—have attracted growing interest in low‑rank optimization. The tangent cone to low-rank sets is widely studied and underpins a range of geometric methods. The second-order geometry, which encodes curvature information, is more intricate. In this work, we develop a unified framework to derive explicit formulas for both first- and second-order tangent sets to various low-rank sets, including low-rank matrices, tensors, symmetric matrices, and positive semidefinite matrices. The framework also accommodates the intersection of a low-rank set and another set satisfying mild assumptions, thereby yielding a tangent intersection rule. Through the lens of tangent sets, we establish a necessary and sufficient condition under which a nonsmooth problem and its smooth parameterization share equivalent second-order stationary points. Moreover, we exploit tangent sets to characterize optimality conditions for low-rank optimization and prove that verifying second-order optimality is NP-hard. In a separate line of analysis, we investigate variational geometry of the graph of the normal cone to matrix varieties, deriving the explicit Bouligand tangent cone, Fréchet and Mordukhovich normal cones to the graph. These results are further applied to develop optimality conditions for low-rank bilevel programs.
@@ 1 Introduction
First-order variational results on $\mathcal{M}_{\leq r}$ luke2013Mordukhovichcone; cason2013iterative; hosseini2019MordukhovichClarke; olikier2022continuity, including the characterization of its tangent and normal cones, have contributed to low-rank optimization by allowing the derivation of first-order optimality conditions and supporting the development of geometric methods schneider2015Lojaconvergence; olikier2023RFDR; olikier2025PGD; olikier2025gauss. Furthermore, imposing an additional constraint $\mathcal{H}$ on low-rank optimization has attracted growing interest in the geometry of $\mathcal{M}_{\leq r}\cap\mathcal{H}$, which, however, was previously treated on a case-by-case basis for different $\mathcal{H}$ in the literature cason2013iterative; tam2017sparsesdp; li2020jotaspectral; li2023normalboundedaffine; yang2025spacedecouple.
Moreover, the second-order geometry of $\mathcal{M}_{\leq r}$ remains unclear, mainly due to two obstacles: 1) second-order analysis in essence captures curvature information bonnans2000perturbationanalysis; gfrerer2017robinsonstability, thereby appearing more involved than the first-order counterpart; 2) the determinantal variety is both nonconvex and nonsmooth olikier2022continuity; levin2023remedy, hindering the direct use of standard analysis tools.
@@ 2.1 Notation
Let $\mathrm{St}(n,p)=\{X\in\mathbb{R}^{n\times p}\mid X^{\top}X=I_{p}\}$ be the Stiefel manifold, $\mathcal{O}(n)=\mathrm{St}(n,n)$ be the orthogonal group, $\mathcal{M}_{s}=\{X\in\mathbb{R}^{m\times n}\mid\operatorname{rank}(X)=s\}$ be the set of fixed-rank matrices, and $\mathrm{Skew}(n)=\{\Omega\in\mathbb{R}^{n\times n}\mid\Omega^{\top}=-\Omega\}$ be the set of skew-symmetric matrices. We use $\operatorname{Diag}(x)$ to denote the diagonal matrix with diagonal entries given by the vector $x$, and $\operatorname{diag}(X)$ to denote the vector extracting the diagonal from a square matrix $X$. Given a smooth manifold $\mathcal{X}$, $\mathrm{T}_{\mathcal{X}}(X)$ denotes the tangent space at $X\in\mathcal{X}$. Given a mapping $F:\mathcal{X}_{1}\to\mathcal{X}_{2}$ between two manifolds, $\mathrm{D}F_{X}:\mathrm{T}_{\mathcal{X}_{1}}(X)\to\mathrm{T}_{\mathcal{X}_{2}}(F(X))$ denotes the differential of $F$ at $X\in\mathcal{X}_{1}$. The standard inner product in an Euclidean space is given by $\langle{X_{1},X_{2}}\rangle:=\mathrm{tr}(X_{1}^{\top}X_{2})$. The distance from $Y$ to $\mathcal{X}$ is defined as $\operatorname{dist}(Y,\mathcal{X})=\inf_{X\in\mathcal{X}}\|X-Y\|$. Let $\mathcal{P}_{\mathcal{X}}$ denote the projection onto the set $\mathcal{X}$. Given a matrix $\in\mathrm{St}(n,r), VV_{\bot}$ is an orthogonal complement of it in the sense of $V\ V_{\bot}]\in\mathcal{O}(n)$. Throughout the paper, whenever the rank of a matrix $X\in\mathbb{R}^{m\times n}$ is explicitly specified—e.g., $\operatorname{rank}(X)=s$ or $X\in\mathcal{M}_{s}$—and an SVD of $X$ is invoked, we use the following conventions: for the compact SVD, $X=U\varSigma V^{\top}$ with $U\in\mathrm{St}(m,s)$, $\varSigma\in\mathbb{R}^{s\times s}$, and $V\in\mathrm{St}(n,s)$; for the full SVD, $X=\bar{U}[\bar{\varSigma}\ 0]\bar{V}^{\top}$ with $\bar{U}\in\mathcal{O}(m)$, $\bar{\varSigma}\in\mathbb{R}^{m\times m}$, and $\bar{V}\in\mathcal{O}(n)$ (when $m\leq n$). Based on the compact SVD, the Moore–Penrose inverse of $X$ is $X^{\dagger}=V\varSigma^{-1}U^{\top}$. Additionally, we use $\sigma_{i}(X)$ to denote the $i$-th largest singular value of $X\in\mathbb{R}^{m\times n}$; when $X$ is symmetric, i.e., $X\in\mathcal{S}(n)$, we use $\lambda_{i}(X)$ to denote its $i$-th largest eigenvalue. Matrices or vectors are denoted by usual roman letters (e.g., $X$, $\eta$), while higher-order tensors are written in boldface (e.g., $\mathbf{X}$, $\bm{\eta}$).
@@ 3 First- and second-order tangent sets
    it is locally Lipschitz, since $|\sigma_{r+1}(X)-\sigma_{r+1}(X+\Delta)|\leq\left\|{\Delta}\right\|_{2}$ for any $\Delta\in\mathbb{R}^{m\times n}$ by Weyl’s inequality weyl1912inequality;
@@
    and according to the truncated SVD, $\sigma_{r+1}$ can control the distance of points to the set $\mathcal{M}_{\leq r}$, i.e., $\operatorname{dist}(X,\mathcal{M}_{\leq r})\leq(\min\{m,n\}-r)^{1/2}\sigma_{r+1}(X)$ for any $X\in\mathbb{R}^{m\times n}$.
@@ 4.2.2 Low-rank symmetric matrices
We briefly explain the above decomposition. In fact, any $X\in\mathcal{S}(n)$ with $\operatorname{rank}(X)\leq r$ must have at least nrn-r consecutive eigenvalues $\lambda_{i}$ equal zero, and by the non-increasing ordering of the $\lambda_{i}$, we conclude that $X\in\mathcal{S}_{j}$ for some $1\leq j\leq r+1$.
@@ 5 Tangent sets bridge optimization landscapes
where the feasible region $\mathcal{M}$ may be nonsmooth and nonconvex, with a typical example being $\mathcal{M}=\mathcal{M}_{\leq r}$. To circumvent the irregularity of $\mathcal{M}$, a common approach is to employ the technique of smooth parameterization levin2023remedy; rebjock2024boundedrank; gao2024desingularizationtensor; levin2025effect; yang2025spacedecouple, which introduces a smooth manifold $\overline{\mathcal{M}}$ to (over)parameterize $\mathcal{M}$, thereby inducing a Riemannian optimization problem; see Fig. 2 for an illustration and see absil2008optimization; boumal2023introduction for more background of Riemannian optimization.
@@ 5.1 Smooth parameterization for optimization problems
Implementing algorithms directly on $\mathcal{M}$ may suffer from the nonsmoothness. To address this, smooth parameterization is introduced levin2023remedy; rebjock2024boundedrank; levin2025effect; yang2025spacedecouple. Specifically, as demonstrated in Fig. 2, let $\mathcal{M}$ and $\overline{\mathcal{M}}$ denote a (possibly nonsmooth) set in $\mathcal{E}$ and a smooth manifold embedded in $\mathcal{W}$, respectively, and let $\phi:\mathcal{W}\to\mathcal{E}$ be a smooth mapping between the two Euclidean spaces such that $\phi(\overline{\mathcal{M}})=\mathcal{M}$. Through the parameterization $(\overline{\mathcal{M}},\phi)$, the original nonsmooth problem (P) can thus be cast as a smooth Riemannian optimization problem (P-M).
@@ 5.2 Equivalence between second-order stationary points
Let $X=\phi(Y)$. In fact, the mappings $\mathrm{\bf L}$ and $\mathrm{\bf Q}$ convey the geometric information encoded in the tangent sets of $\overline{\mathcal{M}}$ to those of $\mathcal{M}$. To see this, we note that for the manifold $\overline{\mathcal{M}}$, any $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ admits a curve $\gamma(t)$ such that $\gamma(0)=Y,\gamma^{\prime}(0)=v$, and thus $\mathrm{\bf L}_{Y}(v)=(\phi\circ\gamma)^{\prime}(0)$, which implies $\operatorname{im}(\mathrm{\bf L}_{Y})\subseteq\mathrm{T}_{\mathcal{M}}(X)$. Moreover, it is revealed from (rockafellar2009variationalanalysis, Proposition 13.13) that given any $u_{v}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v)$ associated with $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, there exists a curve $\beta(t)$ satisfying $\beta(0)=Y$, $\beta^{\prime}(0)=v$, and $\beta^{\prime\prime}(0)=u_{v}$. Hence we have $\mathrm{\bf Q}_{Y,v}(u_{v})=(\phi\circ\beta)^{\prime\prime}(0)$, indicating that $\operatorname{im}(\mathrm{\bf Q}_{Y,v})\subseteq\mathrm{T}^{2}_{\mathcal{M}}(X;\mathrm{\bf L}_{Y}(v))$.
@@
$\nabla\bar{f}(Y)=\mathrm{\bf L}^{*}_{Y}(\nabla f(X))\ \ \text{and}\ \ \nabla^{2}\bar{f}(Y)=\mathrm{\bf L}^{*}_{Y}\circ\nabla^{2}f(X)\circ\mathrm{\bf L}_{Y}+\nabla^{2}\phi_{\nabla f(X)}(Y),$ (5.2)
where $\mathrm{\bf L}^{*}_{Y}:\mathcal{E}\to\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ is the adjoint of $\mathrm{\bf L}_{Y}$, and the mapping $\phi_{\eta}:\mathcal{W}\to\mathbb{R}$ is given by $\phi_{\eta}(W)=\langle{\eta,\phi(W)}\rangle$ for any $(\eta,W)\in(\mathcal{E},\mathcal{W})$. Then, given a pair $(v,u_{v})$, we can specify the following computation,
@@
$=$ $\ \langle{\mathrm{\bf L}_{Y}^{*}(\nabla f(X)),u_{v}}\rangle+\left\langle v,\left({\mathrm{\bf L}^{*}_{Y}\circ\nabla^{2}f(X)\circ\mathrm{\bf L}_{Y}+\nabla^{2}\phi_{\nabla f(X)}(Y)}\right)[v]\right\rangle$
$=$ $\ \langle{\nabla f(X),\mathrm{\bf L}_{Y}(u_{v})}\rangle+\left\langle\mathrm{\bf L}_{Y}(v),\nabla^{2}f(X)[\mathrm{\bf L}_{Y}(v)]\right\rangle+\langle{\nabla f(X),\mathrm{D}^{2}\phi_{Y}[v,v]}\rangle$
@@
where the first equality comes by substituting the derivatives of $\bar{f}$ (5.2). Moreover, we say that a sequence $\{\zeta_{i}+Z\}_{i\in\mathbb{N}}$ of translates of a subspace $Z\subseteq\mathcal{E}$ converges (necessarily to a translate of $Z$) if there exists a sequence $\{z_{i}\}_{i\in\mathbb{N}}\subseteq Z$ such that $\{\zeta_{i}+z_{i}\}_{i\in\mathbb{N}}$ converges.
In the following theorem, we use $\overline{\mathrm{conv}}(\cdot)$ to denote the closed convex hull of a set, and propose a sufficient and necessary condition for the property $2\!\Rightarrow\!2$”. More specifically, when the condition is violated, the failure of $2\!\Rightarrow\!2$” can be witnessed by an explicitly constructed objective $f$.
@@
$\mathrm{T}^{2}_{\mathcal{M}}(X;\mathrm{\bf L}_{Y}(v))\subseteq\overline{\mathrm{conv}}\left({\bigcup_{\{v_{i}\}_{i\in\mathbb{N}}:\,\mathrm{\bf L}_{Y}(v_{i})\to\mathrm{\bf L}_{Y}(v)}\lim_{i\to\infty}\left({\mathrm{\bf Q}_{Y,v_{i}}(u_{v_{i}})+\operatorname{im}(\mathrm{\bf L}_{Y})}\right)}\right).$ (5.4)
@@
We first prove the “if ” part by assuming that $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$ and the inclusion (5.4) holds for all $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. Suppose that $Y$ is second-order stationary for problem (P-M). Then, the first-order condition $\nabla\bar{f}(Y)\in\mathrm{N}_{\overline{\mathcal{M}}}(Y)$ implies that
$0=\langle{v,\nabla\bar{f}(Y)}\rangle=\langle{v,\mathrm{\bf L}_{Y}^{*}(\nabla f(X))}\rangle=\langle{\mathrm{\bf L}_{Y}(v),\nabla f(X)}\rangle,\ \ \text{for all}\ v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y),$

### v2

@@ Abstract
Determinantal varieties—the sets of bounded-rank matrices or tensors—have attracted growing interest in low-rank optimization. The tangent cone to low-rank sets is widely studied and underpins a range of geometric methods. The second-order geometry, which encodes curvature information, is more intricate. In this work, we develop a unified framework to derive explicit formulas for both first- and second-order tangent sets to various low-rank sets, including low-rank matrices, tensors, symmetric matrices, and positive semidefinite matrices. The framework also accommodates the intersection of a low-rank set and another set satisfying mild assumptions, thereby yielding a tangent intersection rule. Through the lens of tangent sets, we establish a necessary and sufficient condition under which a nonsmooth problem and its smooth parameterization share equivalent second-order stationary points. Moreover, we exploit tangent sets to characterize optimality conditions for low-rank optimization and prove that verifying second-order optimality is NP-hard. In a separate line of analysis, we investigate variational geometry of the graph of the normal cone to matrix varieties, deriving the explicit Bouligand tangent cone, Fréchet and Mordukhovich normal cones to the graph. These results are further applied to develop optimality conditions for low-rank bilevel programs.
@@ 1 Introduction
First-order variational results on $\mathcal{M}_{\leq r}$ luke2013Mordukhovichcone; cason2013iterative; hosseini2019MordukhovichClarke; olikier2022continuity, including the characterization of its tangent and normal cones, have contributed to low-rank optimization by allowing the derivation of first-order optimality conditions and supporting the development of geometric methods schneider2015Lojaconvergence; gao2022rankadap; olikier2023RFDR; olikier2025PGD; olikier2025gauss. Furthermore, imposing an additional constraint $\mathcal{H}$ on low-rank optimization has attracted growing interest in the geometry of $\mathcal{M}_{\leq r}\cap\mathcal{H}$, which, however, was previously treated on a case-by-case basis for different $\mathcal{H}$ in the literature cason2013iterative; tam2017sparsesdp; li2020jotaspectral; li2023normalboundedaffine; yang2025spacedecouple.
Moreover, the second-order geometry of $\mathcal{M}_{\leq r}$ remains unclear, mainly due to two obstacles: 1) second-order analysis in essence captures curvature information bonnans2000perturbationanalysis; gfrerer2017robinsonstability, thereby appearing more involved than the first-order counterpart; 2) the determinantal variety is both nonconvex and nonsmooth olikier2022continuity; levin2023remedy, hindering the direct use of standard tools in variational analysis.
@@ 2.1 Notation
Let $\mathrm{St}(n,p)=\{X\in\mathbb{R}^{n\times p}\mid X^{\top}X=I_{p}\}$ be the Stiefel manifold, $\mathcal{O}(n)=\mathrm{St}(n,n)$ be the orthogonal group, $\mathcal{M}_{s}=\{X\in\mathbb{R}^{m\times n}\mid\operatorname{rank}(X)=s\}$ be the set of fixed-rank matrices, and $\mathrm{Skew}(n)=\{\Omega\in\mathbb{R}^{n\times n}\mid\Omega^{\top}=-\Omega\}$ be the set of skew-symmetric matrices. We use $\operatorname{Diag}(x)$ to denote the diagonal matrix with diagonal entries given by the vector $x$, and $\operatorname{diag}(X)$ to denote the vector extracting the diagonal from a square matrix $X$. Given a smooth manifold $\mathcal{X}$, $\mathrm{T}_{\mathcal{X}}(X)$ denotes the tangent space at $X\in\mathcal{X}$. Given a mapping $F:\mathcal{X}_{1}\to\mathcal{X}_{2}$ between two manifolds, $\mathrm{D}F_{X}:\mathrm{T}_{\mathcal{X}_{1}}(X)\to\mathrm{T}_{\mathcal{X}_{2}}(F(X))$ denotes the differential of $F$ at $X\in\mathcal{X}_{1}$. The standard inner product in an Euclidean space is given by $\langle{X_{1},X_{2}}\rangle:=\mathrm{tr}(X_{1}^{\top}X_{2})$. The distance from $Y$ to $\mathcal{X}$ is defined as $\operatorname{dist}(Y,\mathcal{X})=\inf_{X\in\mathcal{X}}\|X-Y\|$. Let $\mathcal{P}_{\mathcal{X}}$ denote the projection onto the set $\mathcal{X}$. Given a matrix $\in\mathrm{St}(n,r), VV_{\bot}$ is an orthogonal complement of it in the sense of $V\ V_{\bot}]\in\mathcal{O}(n)$. Throughout the paper, whenever the rank of a matrix $X\in\mathbb{R}^{m\times n}$ is explicitly specified—e.g., $\operatorname{rank}(X)=s$ or $X\in\mathcal{M}_{s}$—and an SVD of $X$ is invoked, we use the following conventions: for the compact SVD, $X=U\varSigma V^{\top}$ with $U\in\mathrm{St}(m,s)$, $\varSigma\in\mathbb{R}^{s\times s}$, and $V\in\mathrm{St}(n,s)$; for the full SVD, $X=\bar{U}[\bar{\varSigma}\ 0]\bar{V}^{\top}$ with $\bar{U}\in\mathcal{O}(m)$, $\bar{\varSigma}\in\mathbb{R}^{m\times m}$, and $\bar{V}\in\mathcal{O}(n)$ (when $m\leq n$). Based on the compact SVD, the Moore–Penrose inverse of $X$ is $X^{\dagger}=V\varSigma^{-1}U^{\top}$. Additionally, we use $\sigma_{i}(X)$ to denote the $i$-th largest singular value of $X\in\mathbb{R}^{m\times n}$; when $X$ is symmetric, i.e., $X\in\mathcal{S}(n)$, we use $\lambda_{i}(X)$ to denote its $i$-th largest eigenvalue. For convenience, we set $\sigma_{i}\equiv 0$ for all $i>\min\{m,n\}$ and $\lambda_{i}\equiv 0$ for all $i>n$. Matrices or vectors are denoted by usual roman letters (e.g., $X$, $\eta$), while higher-order tensors are written in boldface (e.g., $\mathbf{X}$, $\bm{\eta}$).
@@ 3 First- and second-order tangent sets
    it is locally Lipschitz, since $|\sigma_{r+1}(X)-\sigma_{r+1}(X+\Delta)|\leq\left\|{\Delta}\right\|_{2}$ for any $\Delta\in\mathbb{R}^{m\times n}$ by Weyl’s inequality weyl1912inequality; and
@@
    according to the truncated SVD, $\sigma_{r+1}$ can control the distance of points to the set $\mathcal{M}_{\leq r}$, i.e., $\operatorname{dist}(X,\mathcal{M}_{\leq r})\leq(\min\{m,n\}-r)^{1/2}\sigma_{r+1}(X)$ for any $X\in\mathbb{R}^{m\times n}$.
@@ 4.2.2 Low-rank symmetric matrices
We briefly explain the above decomposition. In fact, any $X\in\mathcal{S}(n)$ with $\operatorname{rank}(X)\leq r$ must have at least nrn-r consecutive eigenvalues $\lambda_{i}$ equal zero, and by the non-increasing ordering of the $\lambda_{i}$, we conclude that $X\in\mathcal{S}_{j}$ for some $1\leq j\leq r+1$, i.e., $\lambda_{1}(X)\geq\cdots\geq\lambda_{j-1}(X)\geq 0=\lambda_{j}(X)=\cdots=\lambda_{j+n-r-1}(X)=0\geq\lambda_{j+n-r}(X)\geq\!\cdots\geq\lambda_{n}(X)$.
@@ 5 Tangent sets bridge optimization landscapes
where the feasible region $\mathcal{M}$ is possibly nonsmooth and nonconvex, with a typical example being $\mathcal{M}=\mathcal{M}_{\leq r}$. To circumvent the irregularity of $\mathcal{M}$, a common approach is to employ the technique of smooth parameterization levin2023remedy; rebjock2024boundedrank; gao2024desingularizationtensor; levin2025effect; yang2025spacedecouple, which introduces a smooth manifold $\overline{\mathcal{M}}$ to (over)parameterize $\mathcal{M}$, thereby inducing a Riemannian optimization problem; see Fig. 2 for an illustration and see absil2008optimization; boumal2023introduction for more background of Riemannian optimization.
@@ 5.1 Smooth parameterization for optimization problems
Implementing algorithms directly on $\mathcal{M}$ may suffer from the nonsmoothness. To address this, smooth parameterization is introduced levin2023remedy; rebjock2024boundedrank; levin2025effect; yang2025spacedecouple. Specifically, as demonstrated in Fig. 2, let $\mathcal{M}$ and $\overline{\mathcal{M}}$ denote a (possibly nonsmooth) set in $\mathcal{E}$ and a smooth manifold embedded in $\mathcal{W}$, respectively, and let $\phi:\mathcal{W}\to\mathcal{E}$ be a smooth mapping between the two Euclidean spaces such that $\phi(\overline{\mathcal{M}})=\mathcal{M}$. Through the parameterization $(\overline{\mathcal{M}},\phi)$, the original nonsmooth problem (P) can thus be cast as a smooth optimization problem on the manifold (P-M).
@@ 5.2 Equivalence between second-order stationary points
Let $X=\phi(Y)$. In fact, the mappings $\mathrm{\bf L}$ and $\mathrm{\bf Q}$ convey the geometric information encoded in the tangent sets of $\overline{\mathcal{M}}$ to those of $\mathcal{M}$. To see this, we note that for the manifold $\overline{\mathcal{M}}$, any $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ admits a curve $\gamma(t)$ such that $\gamma(0)=Y,\gamma^{\prime}(0)=v$, and thus $\mathrm{\bf L}_{Y}(v)=(\phi\circ\gamma)^{\prime}(0)$, which implies $\operatorname{im}(\mathrm{\bf L}_{Y})\subseteq\mathrm{T}_{\mathcal{M}}(X)$. Moreover, it is revealed from (rockafellar2009variationalanalysis, Proposition 13.13) that given any $u_{v}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v)$ associated with $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, there exists a curve $\beta(t)$ satisfying $\beta(0)=Y$, $\beta^{\prime}(0)=v$, and $\beta^{\prime\prime}(0)=u_{v}$. Hence, we have $\mathrm{\bf Q}_{Y,v}(u_{v})=(\phi\circ\beta)^{\prime\prime}(0)$, indicating that $\operatorname{im}(\mathrm{\bf Q}_{Y,v})\subseteq\mathrm{T}^{2}_{\mathcal{M}}(X;\mathrm{\bf L}_{Y}(v))$. The construction of the curve $\beta(t)$ indeed introduces the concept of second fundamental form; see (do1992riemannian, §6.2) and (boumal2023introduction, §5.11) for more details. Specifically, we can define a quadratic form [^3] Different choices of $\beta$ yield equal $\beta^{\prime\prime}(0)$, modulo $\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, ensuring no ambiguity in the definition of $\mathrm{II}_{Y}(v,v)$; see, e.g., (levin2025effect, formula (3.7)). $\mathrm{II}_{Y}(v,v)=\mathcal{P}_{\mathrm{N}_{\overline{\mathcal{M}}}(Y)}(\beta^{\prime\prime}(0))$, for $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, which induces the second fundamental form $\mathrm{II}_{Y}(v,v^{\prime})$:
$\mathrm{II}_{Y}(v,v^{\prime}):=\frac{1}{4}\left[\mathrm{II}_{Y}(v+v^{\prime},v+v^{\prime})-\mathrm{II}_{Y}(v-v^{\prime},v-v^{\prime})\right],$
for $(v,v^{\prime})\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)\times\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. In this view, we have another characterization of the second-order tangent set to the manifold,
$\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v)=\mathrm{II}_{Y}(v,v)+\mathrm{T}_{\overline{\mathcal{M}}}(Y);$
see Lemma 6. Therefore, the image of $\mathrm{\bf Q}_{Y,v}$ is indeed a translate of $\operatorname{im}(\mathrm{\bf L}_{Y})$, i.e.,
$\operatorname{im}(\mathrm{\bf Q}_{Y,v})=\mathrm{D}\phi_{Y}[\mathrm{II}_{Y}(v,v)]+\mathrm{D}^{2}\phi_{Y}[v,v]+\operatorname{im}(\mathrm{\bf L}_{Y}).$ (5.2)
@@
$\nabla\bar{f}(Y)=\mathrm{D}\phi^{*}_{Y}(\nabla f(X))\ \ \text{and}\ \ \nabla^{2}\bar{f}(Y)=\mathrm{D}\phi^{*}_{Y}\circ\nabla^{2}f(X)\circ\mathrm{D}\phi_{Y}+\nabla^{2}\phi_{\nabla f(X)}(Y),$ (5.3)
where $\mathrm{D}\phi^{*}_{Y}:\mathcal{E}\to\mathcal{W}$ is the adjoint of $\mathrm{D}\phi_{Y}$, and the mapping $\phi_{\eta}:\mathcal{W}\to\mathbb{R}$ is given by $\phi_{\eta}(W)=\langle{\eta,\phi(W)}\rangle$ for any $(\eta,W)\in(\mathcal{E},\mathcal{W})$. Then, given $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ and $u_{v}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v)$, we can specify the following computation,
@@
$=$ $\ \langle{\mathrm{D}\phi^{*}_{Y}(\nabla f(X)),u_{v}}\rangle+\left\langle v,\left({\mathrm{D}\phi^{*}_{Y}\circ\nabla^{2}f(X)\circ\mathrm{D}\phi_{Y}+\nabla^{2}\phi_{\nabla f(X)}(Y)}\right)[v]\right\rangle$
$=$ $\ \langle{\nabla f(X),\mathrm{D}\phi_{Y}(u_{v})}\rangle+\left\langle\mathrm{\bf L}_{Y}(v),\nabla^{2}f(X)[\mathrm{\bf L}_{Y}(v)]\right\rangle+\langle{\nabla f(X),\mathrm{D}^{2}\phi_{Y}[v,v]}\rangle$
@@
where the first equality comes by substituting the derivatives of $\bar{f}$ (5.3), and the second equality holds by identifying $\mathrm{D}\phi_{Y}=\mathrm{\bf L}_{Y}$ on $\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. Moreover, we say that a sequence $\{\zeta_{i}+Z\}_{i\in\mathbb{N}}$ of translates of a subspace $Z\subseteq\mathcal{E}$ converges (necessarily to a translate of $Z$) if there exists a sequence $\{z_{i}\}_{i\in\mathbb{N}}\subseteq Z$ such that $\{\zeta_{i}+z_{i}\}_{i\in\mathbb{N}}$ converges.
In Theorem 5.1, We resort to the concept of the following sets:
$\mathcal{Q}_{Y}(v):=\bigcup_{\{v_{i}\}_{i\in\mathbb{N}}:\,\mathrm{\bf L}_{Y}(v_{i})\to\mathrm{\bf L}_{Y}(v)}\lim_{i\to\infty}\left({\mathrm{\bf Q}_{Y,v_{i}}(u_{v_{i}})+\operatorname{im}(\mathrm{\bf L}_{Y})}\right),\ \ \text{for}\ v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y),$
$\varGamma_{Y}:=\left\{{\mathcal{P}_{\hat{\mathrm{N}}_{\mathcal{M}}(\phi(Y))}\left({\mathrm{D}\phi_{Y}[\mathrm{II}(v_{0},v_{1})]+\mathrm{D}^{2}\phi_{Y}[v_{0},v_{1}]}\right)\mid v_{0}\in\operatorname{ker}(\mathrm{\bf L}_{Y}),\,v_{1}\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)}\right\}.$ (5.5)
We use $\overline{\mathrm{conv}}(\cdot)$ to denote the closed convex hull of a set, and propose a sufficient and necessary condition for the property $2\!\Rightarrow\!2$”. More specifically, when the condition is violated, the failure of $2\!\Rightarrow\!2$” can be witnessed by an explicitly constructed objective. The analysis invokes several lemmas capturing the properties of the constructed sets $\mathcal{Q}_{Y}(v)$ and $\varGamma_{Y}$, which are presented in Appendix E.
@@
$\mathrm{T}^{2}_{\mathcal{M}}(X;\mathrm{\bf L}_{Y}(v))\subseteq\overline{\mathrm{conv}}\left({\mathcal{Q}_{Y}(v)+\mathcal{Q}_{Y}(0)+\varGamma_{Y}}\right).$ (5.6)
@@
(Sufficiency) We first assume that $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$ and the inclusion (5.6) holds for all $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. Suppose that $Y$ is second-order stationary for problem (P-M). Then, the first-order condition $\nabla\bar{f}(Y)\in\mathrm{N}_{\overline{\mathcal{M}}}(Y)$ implies that
$0=\langle{v,\nabla\bar{f}(Y)}\rangle=\langle{v,\mathrm{D}\phi_{Y}^{*}(\nabla f(X))}\rangle=\langle{\mathrm{\bf L}_{Y}(v),\nabla f(X)}\rangle,\ \ \text{for all}\ v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y),$